Sheridan Airlines is considering two alternatives for the financing of a purchase of a fleet of airplanes. These two alternatives are:

1. Issue 82,050 shares of common stock at $30 per share. (Cash dividends have not been paid nor is the payment of any contemplated.)
2. Issue 9%, 10-year bonds at face value for $2,461,500.


It is estimated that the company will earn $888,000 before interest and taxes as a result of this purchase. The company has an estimated tax rate of 40% and has 119,000 shares of common stock outstanding prior to the new financing.

Determine the effect on net income and earnings per share for these two methods of financing. (Round earnings per share to 2 decimal places, e.g. 2.25.)
Answer #1

Answer:

Here, it is not clear regarding the date of issue of shares or bonds.

It is assumed that shares or bonds are issued at the time of beginning of the financial year.

Particulars Alternative - 1 Alternative - 2
Net Income before Interest and taxes $888,000 $888,000
Less: Interest Expense                          -   ($221,535)
Net Income before taxes $888,000 $666,465
Less: Tax @ 40% ($355,200) ($266,586)
A Net Income after taxes $532,800 $399,879
B No. of shares outstanding 201,050 119,000
Earnings Per Share (A/B) $2.65 $3.36

Alternative- 1 is issue of shares only

Alternative - 2 is issue of bonds only

Interest Expense for the year under Alternative - 2 = $2,461,500 * 9% = $221,535

Total no.of shares outstanding for Alternative- 1 = 119,000 + 82,050 = 201,050 shares

By issuance of common stock there is no contemplation to pay dividends, hence the net income after taxes under Alternative -1 is higher than Alternative -2.

However issue of more shares certainly reduces the Earnings Per Share (EPS), hence the EPS under Alternative -1 is lower than Alternative -2.